# Markov Perfect Equilibria in

### v1

@@ 1 Introduction
Mean-field games (MFGs) were independently introduced by Lasry and Lions [42, 41, 43] and Huang, Caines and Malhamé [37, 38, 39, 40] as a framework to approximate large symmetric games in which the interaction is through the empirical distribution of agents. Since their introduction, they have been widely applied, ranging from heterogeneous agent models in economics to models of jet-lag recovery in neuroscience, see for example [12, 13, 14] for an overview.
@@
In this work, we are interested in the existence and uniqueness of MPE in discrete-time, finite-state games, both in the finite-player and mean-field settings. Stochastic finite-player games in discrete time are classical in game theory [5], and starting from Gomes, Mohr and Souza [30], the mean-field version has been extensively studied theoretically [47, 24, 8, 29], in the context of learning procedures [49, 33, 46] and as approximations for continuous-time MFGs [1, 34, 29]. Related solution concepts for approximating Markov perfect equilibria in games with a large number of players have been proposed in economic theory, see Adlakha and Johari [2], Weintraub, Benkard and Van Roy [48] and the references therein. In the finite-player (resp. mean-field) case, we characterize symmetric MPE by an equation which we call the finite-player (resp. mean-field) Nash-Lasry-Lions equation ($N$-NLL). In the mean-field limit, this equation is known as the master equation. While classical and various weak solutions to the NLL equation in the continuous-time setting have been studied extensively, see for example [10, 17, 19, 27, 9, 44, 45], NLL equations in discrete time have received less attention.
In the finite-player problem, we show that while solutions to ($N$-NLL) are generally not unique, uniqueness is remarkably restored for all sufficiently small time steps without relying on any monotonicity conditions, underscoring the importance of inertia in dynamic games. Additionally, as the time step goes to zero, we prove that the solution of the discrete-time $N$-NLL equation converges to its continuous-time counterpart. Convergence of time discretizations in the mean-field problem have been well-studied [1, 34, 23, 29].
In the mean-field setting, MPE correspond to solutions of the mean-field Nash-Lasry-Lions equation (MF-NLL). This equation is well-defined as it does not involve derivatives, and any solution induces a solution to the mean-field game for any initial condition. We present an explicit example of a one-step game where uniqueness is lost when the time step is large, and we establish existence and uniqueness in a general game when the time horizon and time step are sufficiently small. In continuous time, it is well-known that the NLL equation may develop singularities in finite time, see [18, 6] for an explicit example, and, hence, classical solutions only exist in short time in the absence of monotonicity conditions [31, 28, 11, 3]. Our result can be seen as the discrete-time analogue of these facts. Additionally, in the short-time regime, we prove that solutions to the finite-player NLL equation converge to their mean-field counterpart in discrete time—analogous to established results in the continuous-time setting, see for example [31].
@@
Notation. For ${\mathcal{X}}:=\{1,\ldots,d\}$, the space $\mathbb{R}^{\mathcal{X}}$ is identified with $\mathbb{R}^{d}$, and we interchangeably use the notation $a(x)=a_{x}$ to denote the $x$th coordinate of any $a\in\mathbb{R}^{d}$. The $(d-1)$-dimensional vector $a_{-x}$ is defined by deleting the $x$th entry from $a$. Further, if $\psi:A\times B\to C$ is a function, then $\psi(a):B\to C$ is defined by $\psi(a)(b):=\psi(a,b)$. For $p\in[0,\infty]$, the $p$-norm is denoted by $|\cdot|_{p}$. For $p=0$, we recall the definition of the 00-norm: $|a|_{0}=|\{x:a_{x}\neq 0\}|$. For a function $\psi:\mathbb{R}^{k}\to\mathbb{R}^{k^{\prime}}$, we define its Lipschitz constant with respect to the $|\cdot|_{1}$-norm by $\mathrm{Lip}_{1}(\psi):=\sup_{\tilde{\mu}\neq\mu}|\psi(\tilde{\mu})-\psi(\mu)|_{1}/|\tilde{\mu}-\mu|_{1}$. Finally, probability distributions/vectors are treated as row vectors.
@@ 2.1 The general finite-player problem
We consider a dynamic stochastic game between $N+1$ players or agents indexed by $\{0,\ldots,N\}$ in discrete time over a finite time horizon $\bar{\mathcal{T}}=\{0,h,2h,\ldots,Kh\}$. Here, $h>0$ is the time step and we denote the time horizon by $T=Kh$. We write ${\mathcal{T}}=\bar{\mathcal{T}}\setminus\{T\}$. Let us fix a player $0\leq i\leq N$. Player $i$’s state is described by a discrete-time stochastic process $(X^{i}_{t})_{t\in\bar{\mathcal{T}}}$ taking values in a finite set ${\mathcal{X}}=\{1,\ldots,d\}$. [^5] The specific choice of ${\mathcal{X}}$ is irrelevant. In particular, we do not assume any ordering on ${\mathcal{X}}$. This player uses a feedback policy $\alpha^{i}:{\mathcal{T}}\times{\mathcal{X}}^{N+1}\to\Sigma^{d-1}$ to control the transition probabilities of $X^{i}_{t}$. The feedback control $\alpha^{i}$ is allowed to depend on the current time and position of the $N+1$ players. Here, the set of probability measures on ${\mathcal{X}}$ is identified with the $(d-1)$-dimensional probability simplex $\Sigma^{d-1}:=\{p\in[0,1]^{d}:\sum_{x}p_{x}=1\}$. Given feedback policies $\alpha=(\alpha^{0},\ldots,\alpha^{(N)})$ and given players’ states $(x^{0},\ldots,x^{N})$ at time $t$, the states of players at time $t+h$ are drawn independently of each other. More precisely, let $X=(X^{0},\ldots,X^{N})$ be the state process which tracks the positions of all players. Without loss of generality, we let $X$ be the canonical element in the space $\Omega=({\mathcal{X}}^{N+1})^{\bar{\mathcal{T}}}$. Its $\mathbb{P}^{\alpha}$-transition probabilities are given by
@@
$\ell^{i}:{\mathcal{X}}^{N+1}\times[0,\infty)^{d}\to[0,\infty)$ and $g^{i}:{\mathcal{X}}^{N+1}\to[0,\infty)$ are running and terminal cost functions, respectively, and we use the notation,
@@
for any $0\leq i,j\leq N$. ${\mathcal{J}}^{i}$ is the cost functional while $v^{i}$ is the value function of player $i$. For future use, for a given strategy profile of all players $\alpha=(\alpha^{0},\ldots,\alpha^{(N)})$, we define the vector of strategies of players other than $i$ by $\alpha^{-i}:=(\alpha^{j})_{j\neq i}$.
@@
We recall some facts about discrete dynamic programming equations in Appendix A.1. Finding MPEs amounts to solving a coupled system of dynamic programming equations. This set of equations is commonly referred to as the Nash system and it reads, for $0\leq i\leq N$ and $x\in{\mathcal{X}}^{N+1}$,
$\left\{\begin{split}v^{i}(t,x)&=\underset{a^{i}\in\Sigma^{d-1}}{\inf}\ \mathscr{H}^{i}(x,v^{i}(t+h),a^{i};\alpha^{-i}(t,x)),\quad 0\leq t\leq T,\\ \alpha^{i}(t,x)&=\underset{a^{i}\in\Sigma^{d-1}}{\mathrm{arg\,min}}\ \mathscr{H}^{i}(x,v^{i}(t+h),a^{i};\alpha^{-i}(t,x)),\quad 0\leq t<T,\\ v^{i}(T,x)&=g^{i}(x).\end{split}\right.$ (Nash system)
@@
In the next section, we consider a game where $(\ell^{i},g^{i})=(\ell,g)$ for all $0\leq i\leq N$, for some common cost functions $\ell,g$, and show how (Nash system) reduces to the Nash-Lasry-Lions equation. We note that the techniques introduced in the next section to prove existence and uniqueness of the Nash-Lasry-Lions equation can also be used to establish existence and uniqueness of the (Nash system).
@@ 2.2 The symmetric finite-player problem
We now consider a game with symmetric players and derive the Nash-Lasry-Lions equation. The assumption of symmetry will allow us to obtain a mean-field game in the limit $N\to\infty$. More precisely, we assume that, for all $0\leq i\leq N$, the cost functions are of mean-field type,
@@
and we identify the set of empirical probability measures on ${\mathcal{X}}$ with $\Sigma^{d-1}_{N}$. The analysis simplifies in this symmetric case. We fix a player and call it the tagged player, and we refer to the untagged players as other players. The state process of the problem can be summarized by a Markov chain $(X_{t},Z_{t})$ taking values in $\mathscr{X}_{N}$. $X_{t}$ now denotes the state of the tagged player at time $t$ while $Z_{t}$ refers to the distribution of other players across states at time $t$. Given a feedback control $\alpha(t,x,z)$ of the tagged player and a feedback control $\beta(t,x,z)$ that is used by every other player, the transition probabilities are given by
@@
where $\gamma^{\beta}(t,x,z)\in\Sigma^{d-1}$ denotes the probability distribution of a random variable $Z^{\prime}(x,z,\beta(t))$, constructed as follows. Here, with an abuse of notation we let $\alpha\otimes\beta$ denote the joint control in which the tagged player uses $\alpha$ and all other players use $\beta$. Fix a time $t\in{\mathcal{T}}$, a position of the tagged player $x\in{\mathcal{X}}$ and an empirical distribution $z\in\Sigma^{d-1}_{N}$ of untagged players. At time $t$ there are $Nz_{y}$ many untagged players in a state $y\in{\mathcal{X}}$. Given that the tagged player is in state $x$, each untagged player in state $y$ chooses a next state independently according to the probability law $\beta(t,y,z+\frac{1}{N}(e_{x}-e_{y}))$ where $e_{x}\in\{0,1\}^{d}$ denotes the $x$th unit vector. It is clear that the distribution of these agents is given by the random variable $Z^{\prime}(x,y,z,\beta(t))$ with
@@
Our setting accommodates models with a minimum transition probability $h\sigma^{2}>0$ by setting $A(h,x)\equiv A(h)$ with $A(h)=\{a\in\Sigma^{d-1}\,:\,a(y)\geq h\sigma^{2}\ \forall y\in{\mathcal{X}}\}$. We note that the upper bound $m=m(d)$ on the allowed transitions may depend on $d$.
@@
$S(x)=\{x-2\pi/d,x,x+2\pi/d\},\quad m=m(d)=3.$
To unify the treatment with the mean-field case, we let running and terminal cost functions be given,
@@
In the same way that the Nash system characterizes Markov perfect equilibria, there is a single equation, which we call the Nash-Lasry-Lions equation, whose solution corresponds to symmetric Markov perfect equilibria. It reads, for $(t,x,z)\in\bar{\mathcal{T}}\times\mathscr{X}_{N}$,
$\left\{\begin{split}v(t,x,z)&=\underset{a\in A(h,x)}{\inf}\,H(x,z,E(x,z,v(t+h),\alpha(t)),a),\quad 0\leq t\leq T,\\ \alpha(t,x,z)&=\underset{a\in A(h,x)}{\mathrm{arg\,min}}\,H(x,z,E(x,z,v(t+h),\alpha(t)),a),\quad 0\leq t<T,\\ v(T,x,z)&=g(x,z),\end{split}\right.$ ($N$-NLL)
@@
the expectation is taken with respect to $Z^{\prime}(x,z,\beta)$ introduced in equation (2.3), and for $(x,z,v,a)\in\mathscr{X}_{N}\times\mathbb{R}^{d}\times\Sigma^{d-1}$, we defined another Hamiltonian by
$H(x,z,v,a)=\ell(x,z,a/h)h+a\cdot v.$ (2.6)
The unknowns of equation ($N$-NLL) are $(v,\alpha)$, and we note that $\alpha(t)$ appears on both sides of the equation in the second line.
The following assumption underlies most of our results in this paper. Recall the notation $\mathscr{X}={\mathcal{X}}\times\Sigma^{d-1}$.
@@
  2. (ii)
    (Unique minimizer). The function $A(h,x)\ni a\mapsto\ell(x,\mu,a/h)h+a\cdot v$ admits a unique minimizer for fixed $h>0$ and $(x,\mu,v)\in\mathscr{X}\times\mathbb{R}^{d}$.
@@
    (Boundedness). The terminal cost $g$ is bounded. Furthermore, for every $(x,\mu,h)\in\mathscr{X}\times(0,\infty)$, there exists an $a^{*}(x,\mu,h)\in A(h,x)$ such that
    $\limsup_{h\downarrow 0}\sup_{(x,\mu)\in\mathscr{X}}|\ell(x,\mu,a^{*}(x,\mu,h)/h)|h<\infty.$
We further observe that the boundedness assumption implies that, for any $\beta\in\mathscr{A}_{N}$, the value function $v^{\beta}$ is uniformly bounded in $(t,x,\mu,h)\in{\mathcal{T}}\times\mathscr{X}\times(0,\bar{h})$ for any finite $\bar{h}<\infty$. If $\ell(x,\mu,\cdot)$ is quadratic in $a_{-x}$ and transition probabilities of the form $\sigma^{2}h$, for some $\sigma^{2}\geq 0$, are admissible, then $a^{*}_{y}(x,\mu,h)=\sigma^{2}h$ for $y\neq x$ satisfies the above assumption.
@@
Under Assumption 2.5, standard results on dynamic programming (see Lemma A.1) imply that a given control $\alpha^{*}\in\mathscr{A}_{N}$ is a symmetric Markov perfect equilibrium if and only if there exists a solution $v:\bar{\mathcal{T}}\times\mathscr{X}_{N}\to[0,\infty)$ to the ($N$-NLL) equation with $\alpha=\alpha^{*}$.
For our uniqueness results, we need the following stronger version of Assumption 2.5 (ii).
@@
for all $(x,\mu,a,a^{\prime})\in\mathscr{X}\times[0,\infty)^{2d}$ and $\phi\in\partial_{a}\ell(x,\mu,a)$ that satisfy $a_{y}=a^{\prime}_{y}=0$ for all $y\notin S(x)$ and $\sum_{y\in{\mathcal{X}}}a_{y}=\sum_{y\in{\mathcal{X}}}a^{\prime}_{y}$ Here, $\partial_{a}\ell(x,\mu,a)$ denotes the subgradient of $\ell(x,\mu,\cdot)$ at the point $a$.
@@
If $m<d$, then this restricts the number of possible transitions at each state. Let $\ell:\mathscr{X}\times\mathbb{R}^{d}\to[0,\infty)$ be a mapping that instead of (2.7) satisfies a standard definition of strong convexity with respect to the $|\cdot|_{2}$-norm,
@@
for all $(a,\tilde{a})\in\mathbb{R}^{2d}$ that satisfy $a_{y}=\tilde{a}_{y}=0$ for any $y\notin S(x)$ and $\sum_{y}\tilde{a}_{y}=\sum_{y}a^{\prime}_{y}$. Note that $\nabla_{a}\ell(x,a)$ is a $d$-dimensional vector whose $x$th coordinate is 00.
@@
    Uniqueness: In addition, let Assumption 2.7 hold. Then there exists $h_{*}(N,T)>0$ such that, for every $h<h_{*}(N,T)$, the solution $v$ to the Nash-Lasry-Lions equation ($N$-NLL) is unique.
@@
For $N\geq 1$ and $h>0$, let $T^{*}(N,h)\in h\mathbb{Z}_{+}\cup\{\infty\}$ denote the maximal time duration such that ($N$-NLL) has a unique solution on ${\mathcal{T}}=\{0,h,\ldots,T^{*}(N,h)\}$. Then $T^{*}(N,h)\to\infty$ as $h\downarrow 0$.

### v2

@@ 1 Introduction
Mean-field games (MFGs) were independently introduced by Lasry and Lions [42, 41, 43] and Huang, Caines and Malhamé [37, 38, 39, 40] as a framework to approximate large symmetric games in which the interaction is through the empirical distribution of agents. Since their introduction, they have been widely applied, ranging from heterogeneous agent models in economics to models of jet-lag recovery in neuroscience, see for example [12, 13, 14].
@@
In this work, we are interested in the existence and uniqueness of MPE in discrete-time, finite-state games, both in the finite-player and mean-field settings. Stochastic finite-player games in discrete time are classical in game theory [5], and starting from Gomes, Mohr and Souza [30], the mean-field version has been extensively studied theoretically [47, 24, 8], in the context of learning procedures [49, 33, 46] and as approximations for continuous-time MFGs [1, 34, 29]. Related solution concepts for approximating Markov perfect equilibria in games with a large number of players have been proposed in economic theory, see Adlakha and Johari [2], Weintraub, Benkard and Van Roy [48] and the references therein. In the finite-player (resp. mean-field) case, we characterize symmetric MPE by an equation which we call the finite-player (resp. mean-field) Nash-Lasry-Lions equation (NLL). In the mean-field limit, this equation is known as the master equation. While classical and various weak solutions to the NLL equation in the continuous-time setting have been studied extensively, see for example [10, 17, 19, 27, 9, 44, 45], NLL equations in discrete time have received less attention.
In the finite-player problem, we show that while solutions to the finite-player NLL equation are generally not unique, uniqueness is remarkably restored for all sufficiently small time steps without relying on any monotonicity conditions. Intuitively, the smallness of the time step renders certain transitions suboptimal, thereby underscoring the importance of inertia in dynamic games. Additionally, as the time step goes to zero, we prove that the solution of the discrete-time $N$-NLL equation converges to its continuous-time counterpart. Convergence of time discretizations in the mean-field problem have been well-studied [1, 34, 23, 29].
In the mean-field setting, MPE correspond to solutions of the mean-field Nash-Lasry-Lions equation. This equation is well-defined as it does not involve derivatives, and any solution induces a solution to the mean-field game for any initial condition. We present an explicit example of a one-step game where uniqueness is lost when the time step is large, and we establish existence and uniqueness in a general game when the time horizon and time step are sufficiently small. In continuous time, it is well-known that the NLL equation may develop singularities in finite time, see [18, 6] for explicit examples, and, hence, classical solutions only exist in short time in the absence of monotonicity conditions [31, 28, 11, 3]. Our result can be seen as the discrete-time analogue of these facts. Additionally, in the short-time regime, we prove that solutions to the finite-player NLL equation converge to their mean-field counterpart in discrete time—analogous to established results in the continuous-time setting, see for example [31].
@@
Notation. For ${\mathcal{X}}:=\{1,\ldots,d\}$, the space $\mathbb{R}^{\mathcal{X}}$ is identified with $\mathbb{R}^{d}$, and we interchangeably use the notation $a(x)$ and $a_{x}$ to denote the $x$th coordinate of any $a\in\mathbb{R}^{d}$. The (d1)(d-1)-dimensional vector $a_{-x}$ is defined by deleting the $x$th entry from $a$. Further, if $\psi:A\times B\to C$ is a function, then $\psi(a):B\to C$ is defined by $\psi(a)(b):=\psi(a,b)$. For $p\in[0,\infty]$, the $p$-norm is denoted by $|\cdot|_{p}$. For $p=0$, we recall the definition of the 0-norm: $|a|_{0}:=|\{x:a_{x}\neq 0\}|$. For a function $\psi:\mathbb{R}^{k}\to\mathbb{R}^{k^{\prime}}$, we define its Lipschitz constant with respect to the $|\cdot|_{1}$-norm by $\mathrm{Lip}_{1}(\psi):=\sup_{\tilde{\mu}\neq\mu}|\psi(\tilde{\mu})-\psi(\mu)|_{1}/|\tilde{\mu}-\mu|_{1}$. Finally, probability distributions/vectors are treated as row vectors.
@@ 2.1 The general finite-player problem
We consider a dynamic stochastic game between $N+1$ players or agents indexed by $\{0,\ldots,N\}$ in discrete time over a finite time horizon $\bar{\mathcal{T}}=\{0,h,2h,\ldots,Kh\}$. Here, $h>0$ is the time step and we denote the time horizon by $T=Kh$. We write ${\mathcal{T}}=\bar{\mathcal{T}}\setminus\{T\}$. Let us fix a player $0\leq i\leq N$. Player $i$’s state is described by a discrete-time stochastic process $(X^{i}_{t})_{t\in\bar{\mathcal{T}}}$ taking values in a finite set ${\mathcal{X}}=\{1,\ldots,d\}$. [^5] The specific choice of $\mathcal{X}$} is irrelevant. In particular, we do not assume any ordering on $\mathcal{X}$}. This player uses a feedback policy $\alpha^{i}:{\mathcal{T}}\times{\mathcal{X}}^{N+1}\to\Sigma^{d-1}$ to control the transition probabilities of $X^{i}_{t}$. The feedback control $\alpha^{i}$ is allowed to depend on the current time and position of the $N+1$ players. Here, the set of probability measures on $\mathcal{X}$} is identified with the (d1)(d-1)-dimensional probability simplex $\Sigma^{d-1}:=\{p\in[0,1]^{d}:\sum_{x}p_{x}=1\}$. Given feedback policies $\alpha=(\alpha^{0},\ldots,\alpha^{N})$ and given players’ states $(x^{0},\ldots,x^{N})$ at time $t$, the states of players at time $t+h$ are drawn independently of each other. More precisely, let $X=(X^{0},\ldots,X^{N})$ be the state process which tracks the positions of all players. Without loss of generality, we let $X$ be the canonical element in the space $\Omega=({\mathcal{X}}^{N+1})^{\bar{\mathcal{T}}}$. Its $\mathbb{P}^{\alpha}$-transition probabilities are given by
@@
$\ell^{i}:{\mathcal{X}}^{N+1}\times[0,\infty)^{d}\to[0,\infty)$ and $g^{i}:{\mathcal{X}}^{N+1}\to[0,\infty)$ are running and terminal cost functions, respectively, and we use the notation
@@
for any $0\leq i,j\leq N$. ${\mathcal{J}}^{i}$ is the cost functional while $v^{i}$ is the value function of player $i$. For future use, for a given strategy profile of all players $\alpha=(\alpha^{0},\ldots,\alpha^{N})$, we define the vector of strategies of players other than $i$ by $\alpha^{-i}:=(\alpha^{j})_{j\neq i}$.
@@
We recall some facts about discrete dynamic programming equations in Appendix A.1. Finding MPEs amounts to solving a coupled system of dynamic programming equations. This set of equations is commonly referred to as the Nash system and it reads, for $0\leq i\leq N$, $0\leq t<T$ and $x\in{\mathcal{X}}^{N+1}$,
$\left\{\begin{split}v^{i}(t,x)&=\underset{a^{i}\in\Sigma^{d-1}}{\inf}\ \mathscr{H}^{i}(x,v^{i}(t+h),a^{i};\alpha^{-i}(t,x)),\\ \alpha^{i}(t,x)&=\underset{a^{i}\in\Sigma^{d-1}}{\mathrm{arg\,min}}\ \mathscr{H}^{i}(x,v^{i}(t+h),a^{i};\alpha^{-i}(t,x)),\\ v^{i}(T,x)&=g^{i}(x).\end{split}\right$. (Nash system)
@@
In the next section, we consider a game where $(\ell^{i},g^{i})=(\ell,g)$ for all $0\leq i\leq N$, with some common cost functions $\ell,g$, and show how (Nash system) reduces to the Nash-Lasry-Lions equation. We note that the techniques introduced in the next section to prove the existence and uniqueness of the solutions to the Nash-Lasry-Lions equation can also be used to establish the same for (Nash system).
@@ 2.2 The symmetric finite-player problem
We now consider a game with symmetric players and derive the Nash-Lasry-Lions equation. The assumption of symmetry will allow us to obtain a mean-field game in the limit $N\to\infty$. More precisely, we assume that, for all $0\leq i\leq N$, the cost functions are of mean-field type:
@@
and we identify the set of empirical probability measures on $\mathcal{X}$} with $\Sigma^{d-1}_{N}$. The analysis simplifies in this symmetric case. We fix a player and call it the tagged player, and we refer to the untagged players as other players. The state process of the problem can be described by a Markov chain $(X_{t},Z_{t})$ taking values in $\mathscr{X}_{N}$. $X_{t}$ now denotes the state of the tagged player at time $t$ while $Z_{t}$ refers to the distribution of other players across states at time $t$. Given a feedback control $\alpha(t,x,z)$ of the tagged player and a feedback control $\beta(t,x,z)$ that is used by every other player, the transition probabilities are given by
@@
where $\gamma^{\beta}(t,x,z)\in\Sigma^{d-1}$ denotes the probability distribution of a random variable $Z^{\prime}(x,z,\beta(t))$, constructed as follows. Here, with an abuse of notation, we let $\alpha\otimes\beta$ denote the joint control in which the tagged player uses $\alpha$ and all other players use $\beta$. Fix a time $t\in{\mathcal{T}}$, a position of the tagged player $x\in{\mathcal{X}}$ and an empirical distribution $z\in\Sigma^{d-1}_{N}$ of untagged players. At time $t$, there are $Nz_{y}$ untagged players in a state $y\in{\mathcal{X}}$. Given that the tagged player is in state $x$, each untagged player in state $y$ chooses a next state independently according to the probability law $\beta(t,y,z+\frac{1}{N}(e_{x}-e_{y}))$, where $e_{x}\in\{0,1\}^{d}$ denotes the $x$th unit vector. It is clear that the distribution of these agents is given by the random variable $Z^{\prime}(x,y,z,\beta(t))$ with
@@
Note that setting accommodates models with a minimum transition probability $h\sigma^{2}>0$ by setting $A(h,x)\equiv A(h)$ with $A(h)=\{a\in\Sigma^{d-1}\,:\,a(y)\geq h\sigma^{2}\ \forall y\in{\mathcal{X}}\}$.
@@
$S(x)=\{x-2\pi/d,x,x+2\pi/d\},\quad m=3.$
Remark 2.4.
In general, when considering a class of models varying with $d$, note that the upper bound $m=m(d)$ on the allowed transitions may depend on $d$. In contrast, in the above Example 2.3, the number $m$ is independent of the dimension $d$.
To unify the treatment with the mean-field case, we let running and terminal cost functions be given:
@@
In the same way that the Nash system characterizes Markov perfect equilibria, there is a single equation, which we call the Nash-Lasry-Lions equation, whose solution corresponds to symmetric Markov perfect equilibria. It reads, for $(t,x,z)\in{\mathcal{T}}\times\mathscr{X}_{N}$,
$\left\{\begin{split}v(t,x,z)&=\underset{a\in A(h,x)}{\inf}\,H(x,z,E(x,z,v(t+h),\alpha(t)),a),\\ \alpha(t,x,z)&=\underset{a\in A(h,x)}{\mathrm{arg\,min}}\,H(x,z,E(x,z,v(t+h),\alpha(t)),a),\\ v(T,x,z)&=g(x,z),\end{split}\right$. (NN-NLL)
@@
the expectation is taken with respect to $Z^{\prime}(x,z,\beta)$ introduced in equation (2.3), and for $(x,\mu,v,a)\in\mathscr{X}\times\mathbb{R}^{d}\times\Sigma^{d-1}$, we defined another Hamiltonian by
$H(x,\mu,v,a)=\ell(x,\mu,a/h)h+a\cdot v.$ (2.6)
The unknowns of ($N$-NLL) are $(v,\alpha)$, and we note that $\alpha(t)$ appears on both sides of the equation in the second line.
The following assumptions underlie most of our results in this paper. Recall the notation $\mathscr{X}={\mathcal{X}}\times\Sigma^{d-1}$.
@@
    (Structural). For every $(x,\mu,a)\in\mathscr{X}\times\mathbb{R}^{d}$, the running cost $\ell(x,\mu,a)$ is independent of $a_{x}$.
  2. (ii)
@@
    (Unique minimizer). The function $A(h,x)\ni a\mapsto H(x,\mu,v,a)$ admits a unique minimizer for fixed $h>0$ and $(x,\mu,v)\in\mathscr{X}\times\mathbb{R}^{d}$.
  4. (iv)
    (Boundedness). The terminal cost $g$ is bounded. Furthermore, for every $(x,\mu,h)\in\mathscr{X}\times(0,\infty)$, there exists an $a^{*}(x,\mu,h)\in A(h,x)$ such that lim sup_h0sup_(x,μ)X |(x,μ,a^*(x,μ,h)/h)| h <.
The structural assumption 2.6 (i) is used to pass to the continuous-time limit. The boundedness assumption 2.6 (iv) implies that, for any $\beta\in\mathscr{A}_{N}$, the value function $v^{\beta}$ is uniformly bounded in $(t,x,\mu,h)\in{\mathcal{T}}\times\mathscr{X}\times(0,\bar{h})$ for any finite $\bar{h}<\infty$. If $\ell(x,\mu,\cdot)$ is quadratic in $a_{-x}$ and transition probabilities of the form $\sigma^{2}h$, for some $\sigma^{2}\geq 0$, are admissible, then $a^{*}_{y}(x,\mu,h)=\sigma^{2}h$ for $y\neq x$ satisfies the above assumption.
@@
Under Assumption 2.6, standard results on dynamic programming (see Lemma A.1) imply that a given control $\alpha^{*}\in\mathscr{A}_{N}$ is a symmetric Markov perfect equilibrium if and only if there exists a solution $v:\bar{\mathcal{T}}\times\mathscr{X}_{N}\to[0,\infty)$ to ($N$-NLL) with $\alpha=\alpha^{*}$.
For our uniqueness results, we need the following stronger version of Assumption 2.6 (iii).
@@
for all $(x,\mu,a,a^{\prime})\in\mathscr{X}\times\mathbb{R}^{2d}$ and $\phi\in\partial_{a}\ell(x,\mu,a)$ that satisfy $a_{y}=a^{\prime}_{y}=0$ for all $y\notin S(x)$ and $\sum_{y\in{\mathcal{X}}}a_{y}=\sum_{y\in{\mathcal{X}}}a^{\prime}_{y}$ . Here, $\partial_{a}\ell(x,\mu,a)$ denotes the subgradient of $\ell(x,\mu,\cdot)$ at the point $a$.
@@
If $m<d$, then this restricts the number of possible transitions at each state. Let $\ell:\mathscr{X}\times\mathbb{R}^{d}\to[0,\infty)$ be a mapping that, instead of (2.7), satisfies the following definition of strong convexity with respect to the $|\cdot|_{2}$-norm:
@@
for all $(a,\tilde{a})\in\mathbb{R}^{2d}$ that satisfy $a_{y}=\tilde{a}_{y}=0$ for any $y\notin S(x)$ and $\sum_{y}\tilde{a}_{y}=\sum_{y}a^{\prime}_{y}$. Note that $\nabla_{a}\ell(x,a)$ is a $d$-dimensional vector whose $x$th coordinate is 0.
@@
    Uniqueness: In addition, let Assumption 2.8 hold. Then there exists an $h_{*}(N,T)>0$ such that, for every $h<h_{*}(N,T)$, the solution $v$ to the Nash-Lasry-Lions equation (NN-NLL) is unique.
@@
Let Assumptions 2.6 and 2.8 hold. For $N\geq 1$ and $h>0$, let $T^{*}(N,h)\in h\mathbb{Z}_{+}\cup\{\infty\}$ denote the maximal time duration such that (NN-NLL) has a unique solution on ${\mathcal{T}}=\{0,h,\ldots,T^{*}(N,h)\}$. Then $T^{*}(N,h)\to\infty$ as $h\downarrow 0$.